The hot resistance of an incandescent lamp is 10 ohms and the rated voltage is 50 V. Find the resistance required to operate the lamp from an 80 V supply.


1
Expert's answer
2022-02-06T11:35:12-0500

Resistance of is 10 ohm for 50 volt supply.

Current =50/10=5 A


For current 5 A and voltage 80 volt

Resistance =80/5= 16 ohm